the drawing shows a version of the loop the loop trick for small car. If car is given an initial speed of 4.0 m/s what is the largest value that the radius can have if the the car is to remain in contact with the circular track at all times


1
Expert's answer
2021-03-30T19:54:46-0400

According to conservation of energy, the car must go upward and the difference between the centripetal acceleration and acceleration due to gravity must be greater than 0:


"a_c-g>0,\\\space\\
a_c=\frac{v^2}{R},\\\space\\
\frac{v^2}{R}-g>0,\\\space\\
R<\frac{v^2}{g}."